# Learning Perceptive Humanoid Locomotion over Challenging Terrain

### v1

@@ I INTRODUCTION
Figure 1: Deployment to outdoor environments. We deployed the model in outdoor challenging terrains. Our controller can successfully traverse a range of terrains, including stair, discrete, rough, gravel, sloping, and deep snow terrains.
@@ IV-B Comparisons between Baselines
  • Performance Superiority: Our method reduces linear velocity tracking error by 27.4% (0.265 $\rightarrow$ 0.182) and enhances terrain difficulty handling by 42.9% (5.941 $\rightarrow$ 8.292) compared to the perception-enhanced PPO baseline
@@ IV-C Sensor Noise Resistance
  • Graceful Degradation: Under extreme noise conditions (200% intensity), our method preserves 74.3% of baseline terrain performance ($\frac{6.524}{8.292}$) versus 34.9% ($\frac{1.452}{4.156}$) for transformer architectures
@@
  • Distillation Necessity: Pure reconstruction objectives prove insufficient, exhibiting 22.7% higher velocity tracking error (0.302 vs 0.231 rad/s) at 100% noise levels compared to our distilled approach
@@ IV-E Real World Comparison
Real-world stair experiments demonstrate our method’s superior handling of perception uncertainty. As shown in Fig. 5, the baseline fails due to overconfident foot placement and compounding state estimation errors, leading to unstable transitions and eventual collapse. In contrast, our system achieves reliable ascent through conservative foothold selection with safety margins and full-body stabilization. The world model’s temporal filtering effectively mitigates real-world sensor artifacts like radar dropouts and reflection noise, enabling robust navigation where conventional methods fail catastrophically.

### v3

@@ I INTRODUCTION
Figure 1: Deployment to outdoor environments. We deployed the model in outdoor challenging terrains. Our controller can successfully traverse a range of terrains, including stair, discrete, rough, gravel, sloping, and deep snow terrains. Videos are available at https://www.youtube.com/watch?v=-47gm15wbYA.
@@ IV-B Comparisons between Baselines
  • Performance Superiority: Our method reduces linear velocity tracking error by 31.3% (0.265 $\rightarrow$ 0.182) and enhances terrain difficulty handling by 42.9% (5.941 $\rightarrow$ 8.292) compared to the perception-enhanced PPO baseline
@@ IV-C Sensor Noise Resistance
  • Graceful Degradation: Under extreme noise conditions (200% intensity), our method preserves 74.3% of baseline terrain performance ($\frac{6.524}{8.292}$) versus 40.4% ($\frac{1.452}{3.596}$) for transformer architectures
@@
  • Distillation Necessity: Pure reconstruction objectives prove insufficient, exhibiting 45.4% higher velocity tracking error (0.439 vs 0.302 rad/s) at 100% noise levels compared to our distilled approach
@@ IV-E Real World Comparison
Real-world stair experiments demonstrate our method’s superior handling of perception uncertainty. As shown in Fig. 5, the PPO baseline fails due to overconfident foot placement and compounding state estimation errors, leading to unstable transitions and eventual collapse. In contrast, our system achieves reliable ascent through conservative foothold selection with safety margins and full-body stabilization. The world model’s temporal filtering effectively mitigates real-world sensor artifacts like radar dropouts and reflection noise, enabling robust navigation where conventional methods fail catastrophically.